# A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability

### v2

@@ 4.4.5. Verification without Disclosure
Verification without disclosure lets verifiers confirm policy compliance without accessing private data. These mechanisms rely on cryptographic proofs (usually NIZK) to attest that hidden data satisfies constraints (e.g., (Narula et al., 2018; Wüst et al., 2022; Cecchetti et al., 2017; Bünz et al., 2020; Ma et al., 2021)). The verifier learns only the truth of the statement. Besides the approaches in Operating limits, here we mention those supporting policies beyond basic balance or transfer correctness check (e.g., (Bünz et al., 2020; Miers et al., 2013; Fauzi et al., 2019; Ma et al., 2021; Baudet et al., 2023; Guan et al., 2022; Beer et al., 2024; Papadoulis et al., 2024; Friolo et al., 2024)). zkLedger (Narula et al., 2018) uses Pedersen commitments enabling aggregate computations. zkLedger supports complex auditing via a map-reduce model: banks generate commitments over filtered and transformed data (e.g., checking for non-zero values or squaring for variance) along with NIZKs proofs of correctness, enabling auditors to compute metrics like means and variances. Platypus (Wüst et al., 2022) defines account states with opaque auxiliary data to support custom policy checks. Solidus (Cecchetti et al., 2017) uses PVORM and ZKPs to publicly verify a bank’s transaction log compliance. DSC (Ma et al., 2021) enforces compliance of encrypted transactions with range and authorization proofs.
@@ 4.5. Summary and Takeaways
Table 2. Summary of key technologies for achieving privacy goals and corresponding example protocols
@@
Privacy-preserving payment systems typically combine multiple cryptographic tools, each with distinct trade-offs. Blind signatures are lightweight but centralized; ring signatures provide scalable anonymity at latency cost; threshold encryption and commitments support auditability but not anonymity; and ZKPs (e.g., zkSNARKs, zkSTARKs) offer strong guarantees, but they are computationally intensive, may require substantial RAM (e.g., (Almashaqbeh and Solomon, 2022; Liang et al., 2025)), and require careful engineering for practical deployment. One of the well-known ZKP-based systems is zk-SNARK, which makes transactions compact for verification but needs a trusted setup that, if mishandled, could compromise the security of the system. Importantly, as shown in Section 4.4, implementing auditability often requires careful design of ledger data structures and interaction protocols to avoid unintended information leaks while still enabling policy verification. This multidimensional design space underscores the need for a holistic, design-oriented approach: effective privacy-preserving digital payment systems must integrate layered technical solutions that balance user privacy with institutional accountability.
@@ 5. Open Challenges and Research Directions
Despite significant advances in privacy-preserving digital payment systems, numerous open challenges remain. These challenges stem from the unique intersection of cryptographic capabilities, monetary policy requirements, and societal expectations of privacy and auditability. We highlight a few directions that merit further investigation.
@@ 5.1. Formal Definitions of Privacy and Auditability
Existing systems use ad-hoc definitions for privacy and auditability. As pointed out in different previous works (e.g., (Genkin et al., 2018; Almashaqbeh and Solomon, 2022)), there exists several theoretical gaps related to the definition of unified privacy models and auditability. Most protocols informally claim anonymity, unlinkability, or confidentiality, but lack formal, composable definitions covering graph obfuscation, timing metadata, or value-hiding under active adversaries. Systems combining multiple primitives rarely have UC (Canetti, 2000) or IITM (Kuesters et al., 2013) proofs, and auditability features (e.g., view keys, trapdoors) are seldom formalized. Frameworks for reasoning about selective disclosure, accountability, and conditional anonymity remain missing.
@@ 5.2. Scalability and Efficiency
Many cryptographic protocols and PETs remain computationally demanding (e.g., homomorphic encryption, ZKPs, and MPC). However, real-world payment systems must support high transaction throughput with low latency and high reliability. To bridge this gap, future research must focus on developing optimized, lightweight, and parallelizable privacy-preserving protocols that can operate efficiently in high-throughput environments. Promising directions include hardware acceleration (e.g., via TEEs) and the exploration of protocol variants that balance privacy guarantees with resource efficiency. ZKPs play a central role in many of the systems analyzed, yet several fundamental challenges remain. General-purpose ZKP schemes (e.g., zk-SNARKs, Bulletproofs) face issues related to trusted setup, proof aggregation, and post-quantum security. For instance, zk-SNARKs typically require a trusted setup, introducing centralization risks and limiting composability. While transparent SNARKs (e.g., Plonky2212121https://github.com/0xPolygonZero/plonky2 and Spartan (Setty, 2019)) and STARKs mitigate these concerns, they often incur higher computational costs. Achieving efficient, setup-free ZKPs without compromising performance remains an open research problem. Finally, no current ZKP scheme offers both strong post-quantum guarantees [^22] ZKPs based on bilinear pairings (e.g., Plonk, Groth) or Pedersen vector commitments (i.e., Bulletproof) assume at least the discrete logarithm problem, hence are subject to quantum threats. Other SNARKs (e.g., Ligero, Orion) and STARKs using Merkle tree and other protocols to build a polynomial commitment are post quantum resistant. and performance comparable to elliptic curve-based constructions, highlighting the need for further advances in quantum-resistant proof systems.
@@ 5.3. Balancing Privacy with Regulatory Compliance
Total anonymity is incompatible with regulatory oversight, making auditable privacy and selective disclosure critical design goals. An open challenge is how to enable systems where honest users maintain strong privacy guarantees, while still allowing authorities to investigate illicit activities under well-defined conditions. Several preliminary approaches explore this trade-off, yet significant work remains to formalize disclosure policies, define the types of information subject to auditing, and clarify the roles and interactions of participating entities. Relevant research directions include ZKPs with selective opening, trapdoor, witness encryption, and anonymity budgeting—each enabling varying levels of conditional transparency. For instance, while fully underexplored, witness encryption holds promise for enabling advanced privacy-preserving policies, such as verifiable time locks or conditional disclosures without trusted third parties (see A.3.5). Additionally, emerging requirements such as time-bound anonymity, jurisdiction-aware confidentiality, and programmable disclosure call for more flexible and context-aware privacy controls. Designing PETs that can support such adaptive and policy-compliant behaviors, while remaining robust against misuse, poses a complex challenge involving technical, legal, and usability considerations.
@@ 5.4. Adaptive Privacy and Role Rotation
Privacy-preserving systems rely on anonymity sets, yet managing them securely and adaptively remains an open challenge, particularly in decentralized or evolving environments. Current schemes, such as fixed-size ring signatures or heuristic decoy selection, degrade as usage patterns shift and are vulnerable to poisoning attacks, where adversaries insert tainted outputs to weaken privacy. Cryptographically sound methods for dynamic, attack-resistant anonymity set construction are still lacking. Beyond anonymity sets, adaptive privacy must address the evolving nature of system participants, especially in regulatory or issuance roles. In CBDCs and regulated payment systems, auditors, compliance entities, or issuers may change due to policy or organizational shifts. Supporting this evolution without compromising anonymity or confidentiality is difficult: most systems assume static trust anchors or fixed multi-signature sets. Cryptographic tools such as threshold cryptography, distributed key generation, or policy-based encryption could enable dynamic reconfiguration, but balancing adaptability with unlinkability, correctness, and forward secrecy remains unresolved.
@@ 5.5. Network-Level Anonymity and Cross-layer Privacy Leakage
Network-layer anonymity in crypto-asset systems is typically outsourced to external systems (e.g., Tor, mix networks). However, native integration of network-level privacy mechanisms into privacy-preserving payment systems protocols remains largely unresolved. This delegation creates a critical cross-layer vulnerability, where even perfectly private on-chain transactions can be deanonymized through timing analysis, transaction ordering, volume inference, or network topology correlation. Joint modeling of privacy across the ledger layer and network layer is underexplored. Existing protocols often assume independence between transaction content and communication patterns, which adversaries can exploit. For instance, gossip-based mempool propagation leaks significant metadata about the origin and timing of transactions, enabling traffic correlation attacks. Emerging research directions include the design of privacy-preserving mempools that incorporate indistinguishable broadcasting, cover traffic, oblivious relaying, or dandelion-style (Fanti et al., 2018) diffusion to mask transaction origins. These approaches aim to reduce the information available to global or local adversaries observing network traffic (e.g., (Diaz et al., 2021; Bebel and Ojha, 2022)). However, these solutions often introduce latency, bandwidth overhead, or synchronization issues, leading to difficult trade-offs between efficiency, liveness, and anonymity guarantees. Addressing network anonymity in digital payment systems requires cross-layer designs that align messaging protocols with the cryptographic guarantees of the ledger. This is a highly active area of research, particularly in adversarial environments where well-funded surveillance actors may target both layers simultaneously.
@@ 6. Conclusion
This survey provides a structured and technically grounded overview of the evolving landscape of privacy-preserving digital payment systems. We systematize the literature across untraceable payments, decentralized ledger-based solutions, and CBDCs, introducing a comprehensive taxonomy that connects privacy and auditability goals with underlying cryptographic primitives, protocol designs, and system architectures. In doing so, we bridge conceptual objectives—such as anonymity, confidentiality, and auditability—with concrete technical mechanisms and design models. Our analysis traces the evolution of privacy-preserving digital payment systems across three generations, reflecting an increasing interest in addressing privacy-accountability trade-offs. We underscore that privacy is not a monolithic property but a multifaceted design goal shaped by system requirements and regulatory constraints. Finally, we outline several open research challenges that call for interdisciplinary collaboration across cryptography, distributed systems, economics, and policy. Key directions include the development of scalable and composable privacy primitives, architectures that balance auditability with confidentiality, and formal models that capture realistic adversarial and compliance scenarios. As digital payment technologies progress from academic prototypes to deployed systems, balancing strong privacy guarantees with regulatory compliance will require ongoing technical advances and interdisciplinary efforts. This survey provides a unified taxonomy to guide future research and the design of privacy-preserving payment systems of the future.
@@ Appendix A Appendix
This appendix provides supplementary material for A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability, which complements the main analysis but is excluded from the core sections to preserve narrative flow. Specifically, it includes: (i) a breakdown of publication venues for the works reviewed in this survey, and (ii) a summary of mixing protocols referenced in Section 4.3.5. These materials are presented as tables and offer additional context and technical detail for readers interested in the broader research landscape and protocol-level design trade-offs. Furthermore, in A.3, we briefly present interesting yet underexplored cryptographic tools that can be used to design privacy-preserving digital payment systems.
A.1. Most Popular Publication Venues
Table 3 summarizes the most common publication venues among the reviewed works. We report only the most frequent venues (with more than $2$ works), including journals (e.g., IEEE Transactions on Dependable and Secure Computing, Future Generation Computer Systems), international conferences (e.g., Financial Cryptography and Data Security (FC), ACM Conference on Computer and Communications Security (CCS), IEEE Symposium on Security and Privacy (S&P)), and other (non-peer reviewed) sources, such as preprints and whitepapers.
@@ A.1. Most Popular Publication Venues
Journal 13
IEEE Trans. Dependable Secur. Comput (5)
@@
Conference 30
@@
A.2. Reviewed Mixing Protocols
Similar to Table LABEL:tab:sota-cryptocurrencies, which surveys key privacy-preserving digital payment and crypto-assets systems, Table LABEL:tab:sota-mixing focuses specifically on mixing protocols. The corresponding legend in Table LABEL:tab:sota_mixing_legend extends the notation introduced in Table 1. These mixing protocols are discussed in Section 4.3.5, and the table serves as a reference to facilitate comparison across existing approaches in the literature.

### v3

@@ 4.4.5. Verification without Disclosure
Verification without disclosure lets verifiers confirm policy compliance without accessing private data. These mechanisms rely on cryptographic proofs (usually NIZK) to attest that hidden data satisfies constraints (e.g., (Narula et al., 2018; Wüst et al., 2022; Cecchetti et al., 2017; Bünz et al., 2020; Ma et al., 2021)). The verifier learns only the truth of the statement. Besides the approaches in Operating limits, here we mention those supporting policies beyond basic balance or transfer correctness check (e.g., (Bünz et al., 2020; Miers et al., 2013; Fauzi et al., 2019; Ma et al., 2021; Baudet et al., 2023; Guan et al., 2022; Beer et al., 2024; Papadoulis et al., 2024; Friolo et al., 2026)). zkLedger (Narula et al., 2018) uses Pedersen commitments enabling aggregate computations. It employs generalized Schnorr-based NIZKs to prove the correctness of computations without revealing underlying data. zkLedger supports complex auditing via a map-reduce model: banks generate commitments over filtered and transformed data (e.g., checking for non-zero values or squaring for variance) along with NIZKs proofs of correctness, enabling auditors to compute metrics like means and variances. [^29] Recently, Yang et al. (Yang et al., 2025) extend the zkLedger lineage to permissioned enterprise settings with auditable confidential asset transfers. Platypus (Wüst et al., 2022) defines account states with opaque auxiliary data to support custom policy checks. Solidus (Cecchetti et al., 2017) uses PVORM and ZKPs to publicly verify a bank’s transaction log compliance. DSC (Ma et al., 2021) enforces compliance of encrypted transactions with range and authorization proofs, without revealing transaction details. These approaches primarily support compliance verification rather than attribution: they enable auditors to assess whether rules were followed, without necessarily identifying the transacting parties.
4.4.6. Discussion
Choosing an auditability mechanism depends on the application context and threat model. Lightweight approaches such as anonymity budgets and operating limits are suitable for retail payments where limiting large-scale abuse is sufficient, but they offer limited forensic capabilities and are vulnerable to Sybil strategies unless combined with identity binding. Revocable anonymity provides stronger deterrence by enabling authorized identity recovery, at the cost of increased trust assumptions and governance complexity. Coin-tracing mechanisms support post-incident analysis of fund flows and recovery of tainted assets, but may weaken unlinkability and expose transaction graph structure. Verification without disclosure enables compliance checks (e.g., balance or limit enforcement) without revealing identities or values, but typically does not support attribution. Consequently, no single approach dominates the design space: compliance-oriented systems favor revocable or traceable designs, privacy-first systems rely on limited or aggregate auditability, and hybrid architectures increasingly combine multiple techniques to balance privacy, trust, and enforcement.
@@ 4.5. Summary and Takeaways
Table 2. Summary of privacy goals, main enabling technologies, and representative protocols
@@
Privacy-preserving payment systems typically combine multiple cryptographic tools, each with distinct trade-offs. Blind signatures are lightweight but centralized; ring signatures provide scalable anonymity at latency cost; threshold encryption and commitments support auditability but not anonymity; and ZKPs (e.g., zkSNARKs, zkSTARKs) offer strong guarantees, but they are computationally intensive, may require substantial RAM (e.g., (Almashaqbeh and Solomon, 2022; Liang et al., 2025)), and require careful engineering to achieve practical deployment. One of the well-known ZKP-based systems is zk-SNARK, which makes transactions compact for verification but needs a trusted setup that, if mishandled, could compromise the security of the system. [^30] In practice, parameters are generated via publicly auditable multi-party computation ceremonies, and security holds as long as at least one participant is honest and destroys the corresponding secret randomness. Importantly, as shown in Section 4.4, implementing auditability often requires careful design of ledger data structures and interaction protocols to avoid unintended information leaks while still enabling policy verification. This multidimensional design space highlights the need for a holistic, design-oriented perspective, where privacy-preserving digital payment systems combine complementary mechanisms to balance user privacy with institutional accountability.
@@ 5. Open Challenges and Research Directions
Despite significant advances, numerous open challenges remain at the intersection of business requirements, system design, and expectations of privacy and auditability. We identify a few interesting research directions.
@@ 5.1. Formal Definitions of Privacy and Auditability
Existing systems use ad-hoc definitions for privacy and auditability, and previous works (e.g., (Genkin et al., 2018; Almashaqbeh and Solomon, 2022)) highlight theoretical gaps in unified privacy and auditability models. Most protocols informally claim anonymity, unlinkability, or confidentiality, but lack formal, composable definitions covering graph obfuscation, timing metadata, value-hiding, selective disclosure, and conditional anonymity. Systems combining multiple primitives rarely have UC (Canetti, 2000) or IITM (Küsters et al., 2020) proofs, and auditability features (e.g., view keys, trapdoors) are seldom formalized.
@@ 5.2. Scalability and Efficiency
Many cryptographic protocols and PETs remain computationally demanding (e.g., homomorphic encryption, ZKPs, and MPC). Future research must focus on developing optimized, lightweight, and parallelizable privacy-preserving protocols that can operate efficiently in high-throughput environments. Promising directions include hardware acceleration (e.g., via TEEs) and the exploration of protocol variants that balance privacy guarantees with resource efficiency. ZKPs play a central role in many systems, yet general-purpose schemes (e.g., zk-SNARKs, zk-STARKs) still face challenges related to trusted setup, proof aggregation, and post-quantum security. For instance, zk-SNARKs typically require a trusted setup, introducing centralization risks and limiting composability. While transparent SNARKs (e.g., Plonky2313131https://github.com/0xPolygonZero/plonky2 and Spartan (Setty, 2019)) and STARKs mitigate these concerns, they often incur higher computational costs. Achieving efficient, setup-free ZKPs without compromising performance remains an open research problem. Additionally, support for efficiently generating proofs that can verify other proofs—known as recursive proof composition—is still limited, even though this capability is crucial for scalable applications like off-chain rollups. Developing ZKPs that are not only succinct and recursive, but also universally verifiable (i.e., verifiable by anyone without special access), remains a technically challenging and active area of research. No current ZKP scheme offers simultaneously strong post-quantum guarantees [^32] ZKPs based on bilinear pairings (e.g., Plonk, Groth) or Pedersen vector commitments (i.e., Bulletproof) assume at least the discrete logarithm problem, hence are subject to quantum threats. Other SNARKs (e.g., Ligero, Orion) and STARKs using Merkle tree and other protocols to build a polynomial commitment are post quantum resistant. and performance comparable to elliptic curve-based constructions.
@@ 5.3. Balancing Privacy with Regulatory Compliance
Total anonymity is incompatible with regulatory oversight, making auditable privacy and selective disclosure critical design goals. An open challenge is how to enable systems where honest users maintain strong privacy guarantees, while still allowing authorities to investigate illicit activities under well-defined conditions. While preliminary approaches exist, significant work remains to formalize disclosure policies, define the types of information subject to auditing, and clarify the roles and interactions of participating entities. Relevant research directions include ZKPs with selective opening, trapdoor, witness encryption, and anonymity budgeting—each enabling varying levels of conditional transparency. Less explored primitives, such as witness encryption, may enable advanced privacy-preserving policies (see Appendix A.4.5). Additionally, emerging requirements, such as time-bound anonymity, jurisdiction-aware confidentiality, and programmable disclosure, call for flexible and context-aware privacy controls. Designing PETs that can support such adaptive behaviors, while remaining robust against misuse, poses a complex challenge involving technical, legal, and usability considerations.
@@ 5.4. Adaptive Privacy and Role Rotation
Privacy-preserving systems rely on anonymity sets, yet managing cryptographically sound methods for dynamic and attack-resistant management remain lacking, especially in decentralized or evolving environments. Current schemes, such as fixed-size ring signatures or heuristic decoy selection, degrade as usage patterns shift and are vulnerable to poisoning attacks, where adversaries insert tainted outputs to weaken privacy. Beyond anonymity sets, adaptive privacy must address the evolving nature of system participants, especially in regulatory or issuance roles. In CBDCs and regulated payment systems, auditors, compliance entities, or issuers may change due to policy or organizational shifts. Supporting this evolution without compromising anonymity or confidentiality is difficult: most systems assume static trust anchors or fixed multi-signature sets. Cryptographic tools such as threshold cryptography, distributed key generation, or policy-based encryption could enable dynamic reconfiguration, but balancing adaptability with unlinkability, correctness, and forward secrecy remains unresolved.
@@ 5.5. Network-Level Anonymity and Cross-layer Privacy Leakage
Network-layer anonymity is typically outsourced to external systems (e.g., Tor, mix networks). However, native integration of network-level privacy mechanisms into privacy-preserving payment system protocols remains largely unresolved. This delegation creates a critical cross-layer vulnerability, where even perfectly private on-chain transactions can be de-anonymized through timing analysis, transaction ordering, volume inference, or network topology correlation. Existing protocols often assume independence between transaction content and communication patterns, which adversaries can exploit. For instance, gossip-based mempool propagation leaks significant metadata about the origin and timing of transactions, enabling traffic correlation attacks. Emerging research directions include the design of privacy-preserving mempools that incorporate indistinguishable broadcasting, cover traffic, oblivious relaying, or dandelion-style (Fanti et al., 2018) diffusion to mask transaction origins. These approaches aim to reduce the information available to global or local adversaries observing network traffic (e.g., (Diaz et al., 2021; Bebel and Ojha, 2022)). However, these solutions often introduce latency, bandwidth overhead, or synchronization issues, leading to difficult trade-offs between efficiency, liveness, and anonymity guarantees. Addressing network anonymity in digital payment systems requires cross-layer designs that align messaging protocols with the cryptographic guarantees of the ledger. This remains a highly active research area in adversarial environments.
@@ 6. Conclusion
This survey provides a structured and technically grounded overview of the evolving landscape of privacy-preserving digital payment systems. We systematize the literature across both decentralized and centrally managed designs, including crypto-asset systems and CBDC proposals, introducing a comprehensive taxonomy that connects privacy and auditability goals with underlying cryptographic primitives, protocol designs, and system architectures. In doing so, we bridge conceptual objectives (such as anonymity, confidentiality, and auditability) with concrete technical mechanisms and design models.
Our analysis traces the evolution of privacy-preserving digital payment systems across three generations, reflecting an increasing interest in addressing privacy-accountability trade-offs. We underscore that privacy is not a monolithic property but a multifaceted design goal shaped by system requirements and regulatory constraints. Finally, we outline several open research challenges that call for interdisciplinary collaboration across cryptography, distributed systems, economics, and policy. Key directions include the development of scalable and composable privacy primitives, architectures that balance auditability with confidentiality, and formal models that capture realistic adversarial and compliance scenarios. As digital payment technologies progress from academic prototypes to deployed systems, balancing strong privacy guarantees with regulatory compliance will require ongoing technical advances and interdisciplinary efforts.
@@ Appendix A Appendix
This appendix provides supplementary material for A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability, which complements the main analysis but is excluded from the core sections to preserve narrative flow. Specifically, it includes: (i) the research methodology used for the literature review (Section A.1); (ii) a breakdown of publication venues for the works reviewed in this survey (Section A.2), and (iii) an analysis of mixing protocols (Section A.3) that complements the analysis in Section 4, with particular emphasis on unlinkability techniques discussed in Section 4.3.5. These materials are presented as tables and offer additional context and technical detail for readers interested in the broader research landscape and protocol-level design trade-offs. Furthermore, in Section A.4, we briefly present interesting yet underexplored cryptographic tools that can be used to design privacy-preserving digital payment systems.
A.1. Selection and Classification of the Reviewed Contributions
To identify the relevant contributions we analyzed in the paper, we adopted the following approach:
  1. (1)
    We identified a core collection of works relying on direct authors’ knowledge of the research field;
  2. (2)
    We enlarged this collection through various iterations of backward snowballing (Wohlin, 2014) (i.e., using the references), which allowed us to identify related works not included in the first step;
  3. (3)
    Analyzing the selected papers, we identified the most frequent publication venues (i.e., journals or conferences with 2+ selected papers). Checking the list of papers published in these venues, we added further papers to the collection;
  4. (4)
    As a final step, we performed keyword-based searches on publication databases (i.e., Google Scholar, DBLP, Elsevier Scopus) using combinations of terms such as privacy, auditability/accountability, cryptocurrency/blockchain, crypto-assets, payment systems/CBDC, and specific mechanism keywords (e.g., zero-knowledge proofs, anonymous/confidential transactions, conditional/revocable anonymity, tracing/AML compliance).
After searching and scraping, we selected and reviewed about $70$ publications for this survey. Section A.2 provides information about the most popular publication venues. Table LABEL:tab:sota-cryptocurrencies summarizes the analysis of work related to privacy preserving digital payment systems and crypto-asset systems, whereas Table LABEL:tab:sota-mixing reports the analyzed mixing protocols.
A.2. Most Popular Publication Venues
Table 3 summarizes the most common publication venues among the reviewed works. We report only the most frequent venues (with $2+$ works), including journals (e.g., IEEE Transactions on Dependable and Secure Computing, Future Generation Computer Systems, IEEE Transactions on Information Forensics and Security), international conferences (e.g., Financial Cryptography and Data Security (FC), ACM Conference on Computer and Communications Security (CCS), IEEE Symposium on Security and Privacy (S&P)), and other (non-peer reviewed) sources, such as pre-prints and whitepapers.
@@ A.2. Most Popular Publication Venues
Journal 16
IEEE Trans. Dependable Secur. Comput. (6)
@@
IEEE Trans. Inf. Forensics Secur. (2)
Conference 33
@@
A.3. Analysis of Mixing Protocols
Similar to Table LABEL:tab:sota-cryptocurrencies, which surveys privacy-preserving digital payment and crypto-assets systems, Table LABEL:tab:sota-mixing focuses specifically on mixing protocols. The table serves as a reference to facilitate comparison across existing approaches in the literature. The corresponding symbols in Table LABEL:tab:sota_mixing_legend extend those introduced in Table 1.